OpenStudy (anonymous):

Determine if the lines that pass through the given points are parallel, perpendicular or neither. Line A: (-3, 5) and (0, 7) Line B: (6, 2) and (9, 4)

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

Do you know the slope formula?

OpenStudy (anonymous):

not sure

jimthompson5910 (jim_thompson5910):

hopefully you've seen this formula before \[\Large m = \frac{y_{2}-y_{1}}{x_{2}-x_{1}}\]

jimthompson5910 (jim_thompson5910):

or no?

OpenStudy (anonymous):

oh yes!

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@misty1212

jimthompson5910 (jim_thompson5910):

so you will use that formula to find the slope of the line AB

jimthompson5910 (jim_thompson5910):

do you know how to do that?

OpenStudy (anonymous):

plug in the numbers x1 etc... but i dot have my calc

jimthompson5910 (jim_thompson5910):

you can use online calculators, such as this one http://web2.0calc.com/

jimthompson5910 (jim_thompson5910):

ideally, you should be able to do this without a calculator (just use pen/pencil and paper)

OpenStudy (anonymous):

wait which set do i use

jimthompson5910 (jim_thompson5910):

Slope of AB \[\Large m = \frac{y_{2}-y_{1}}{x_{2}-x_{1}} = \frac{7-5}{0-(-3)} = ??\]

jimthompson5910 (jim_thompson5910):

for AB, you use points A and B you can make A be (x1,y1) and make B be (x2,y2)

OpenStudy (anonymous):

2/3

jimthompson5910 (jim_thompson5910):

yes, slope of AB is 2/3

jimthompson5910 (jim_thompson5910):

oh wait, my bad, this is just the slope of line A

jimthompson5910 (jim_thompson5910):

A is a line, not a point

jimthompson5910 (jim_thompson5910):

slope of line A is 2/3

jimthompson5910 (jim_thompson5910):

Line B: (6, 2) and (9, 4) find the slope of line B

OpenStudy (anonymous):

2/3?

jimthompson5910 (jim_thompson5910):

also, 2/3, yes

jimthompson5910 (jim_thompson5910):

whenever you have equal slopes, you have parallel lines

OpenStudy (anonymous):

so what would the answer be

jimthompson5910 (jim_thompson5910):

extra info: if the two slopes multiply to -1, then the lines are perpendicular.

jimthompson5910 (jim_thompson5910):

it's practically given to you at this point

jimthompson5910 (jim_thompson5910):

slope of line A is 2/3 slope of line B is 2/3

OpenStudy (anonymous):

would it just be 2/3?

jimthompson5910 (jim_thompson5910):

Original Problem: Determine if the lines that pass through the given points are parallel, perpendicular or neither.

OpenStudy (anonymous):

parallel?

jimthompson5910 (jim_thompson5910):

yes, we have 2 equal slopes (both equal to 2/3)

OpenStudy (anonymous):

thank u sososos much